# HET/HPF observations of Helium in warm, hot, and ultra-hot Jupiters

### v1

@@ 2.3 Transmission spectrum analysis
We analyzed the HPF spectroscopic time series using the single line transmission spectroscopy technique, which compares the in-transit spectra with the out-of-transit spectra. We adapted the methodology presented in J. Orell-Miquel et al. (2022, 2023, 2024) to the particularities of the HET following Z. Zhang et al. (2023) and M. Gully-Santiago et al. (2024). Table 3 shows the planetary and stellar parameters needed to compute the transmission spectrum for our targets.
@@
Once corrected from the telluric contributions, we normalized each spectrum by its own continuum. We used muler to fit a first order Chebyshev’s polynomial to featureless spectral regions. Then, we shifted the normalized spectra into the stellar rest frame. We computed a high S/N stellar spectrum from the out-of-transit data (MasterOut). To minimize the impact of stellar variability, we analyzed each in-transit observation individually. For each transit visit, we computed a different MasterOut spectrum using only the out-of-transit spectra that are closer to the in-transit track. We divided all spectra (in-, out-of-transit, and P4) by the MasterOut to remove the stellar contribution and to obtain the residual map in the stellar rest frame. At this step, we computed the stellar light curve of the He i triplet integrating the counts within a 1.5 Å band pass centered at the doublet position. The light curves are shown in the Figures 13 and 14, and provide information of the stellar variability and the presence of large planetary tails (Z. Zhang et al., 2023; M. Gully-Santiago et al., 2024).
@@
We moved the residual map into the planet rest frame accounting for the planet movement around the star. We neglected the planets’ eccentricity because it is usually poorly constrained and the impact is within the uncertainties derived for the signal shifts, except for HD118203 b, HAT-P-17 b, and HAT-P-26 b. For these three planets, we used the radvel (B. J. Fulton et al., 2018) package to compute the radial velocity contribution of the eccentricity during the transit. We computed the transmission spectrum (TS) by averaging the in-transit spectra weighted by the inverse of the squared propagated errors. If we obtained more than one in-transit visit, we computed the final TS from the combined residual maps. When the TS shows an absorption feature, we fitted the signal with a Gaussian profile using the Markov’s chain Monte Carlo algorithm via its python implementation emcee (D. Foreman-Mackey et al., 2013). If there were no clear signals in the TS, we derived conservative limits for the planetary absorption. We computed a 3 upper limit as three times the root-mean-squared value of a flat spectral region of the TS.
3 Transmission spectroscopy results
@@
Figures 9–12 show the performance of the telluric correction and the comparison between the MasterOut and the individual in-transit spectra. For multiple visits, we also present the individual TS (see Figs 10, 11, and 12). The He i triplet stellar light curves are shown in Figures 13 and 14.
@@
3.1 HAT-P-3 b
The He i triplet stellar line partially overlaps with the OH lines in the only visit to the target, but the telluric emission correction is well performed (Fig. 9). The residual map and TS do not indicate the presence of a planetary absorption in our data. We derive a 3 upper limit to the He i excess absorption of 1.2 %.
3.2 HAT-P-12 b
The planetary trace is slightly overlapped by OH emission lines in the two visits (Fig. 10). While the TS from the first visit shows a broad 2 % excess absorption feature, the TS from the second visit is dominated by noise without a clear absorption signal. When combining both observations, final TS (Fig. 2) looks similar to that of the first visit. The shape of the absorption feature does not appear to be of planetary origin, but rather from residual OH emissions. Although, the S/N of the final TS would be sensitive to a He i triplet excess absorption of 1.3 %, we report the 2 % absorption from the signal at the He i triplet position as upper limit.
3.3 HAT-P-17 b
The stellar line does not overlap with the telluric lines, but the planetary trace is in between the OH emission line and the H2O absorption line due to the RV shift caused by the planet’s eccentricity (Fig. 10). We obtained two visits which sampled the terminators of HAT-P-17 b (visit 1: ingress, visit 2: egress). Although we did not get full in-transit spectra, these observations would still be able to detect evaporation tails. Despite the telluric contamination, the TS are mainly flat without significant features (Fig. 10). The combined TS (Fig. 2) does not present absorption features from the planet atmosphere. We derive a 3 upper limit to the He i excess absorption of 0.7 %.
3.4 HAT-P-26 b
The He i stellar line is free from tellurics, but the planetary trace falls over the blue doublet of the OH emission lines when accounting for the RV shift caused by the planet’s eccentricity (Fig. 9). Despite the telluric overlapping, the TS is mainly flat (Fig. 2) and does not suggest He i excess absorption during the transit. The 3 upper limit to the He i excess absorption in our observations is of 1.2 %.
3.5 HD118203 b
@@
3.6 HD189733 b
We obtained three visits that sample almost all the planetary transit (visits 1, 2, and 3 in Fig. 13). The planetary trace does not overlap with the telluric lines (Fig. 9), but the out-of-transit baseline (visit 4 in Fig. 13) is more than one month apart from the in-transit visits. The He i stellar line shows some variability that prevents us from computing the planetary TS. Fortunately, two spectra from P4 observations were taken just a few days after visit 1. We computed the residual map and TS shown in Fig. 3 using those two P4 spectra as out-of-transit and only visit 1 as in-transit. We masked some strong telluric residuals. The TS show a significant absorption feature at the He i triplet position (0.410.08 %). However, we emphasize that there are other features in the TS of the same amplitude.
3.7 HD209458 b
@@ 3.10 WASP-52 b
The first visit is slightly contaminated by OH emission, but the TS presents an absorption at the He i triplet position (Fig. 11). We fitted the feature including an offset since the TS around the He i line is consistently below 0 (offset = 0.550.13 %). We derive a blue-shifted (=73 km s-1) excess absorption of 1.70.4 %. The TS of the second visit shows a clear absorption feature as well. In the second case, we fitted an excess absorption of 1.90.3 % (=13 km s-1). Figure 4 shows the residual map and the TS when combining both visits. Our observations provide a clear He i triplet detection with an excess absorption of 1.850.20 %.
3.11 WASP-69 b
@@
The stellar line from both visits slightly overlap with the OH emission line, but there are no obvious telluric residuals. Although the first visit only has one in-transit spectrum with a useful S/N, the He line is clearly deeper than the MasterOut spectrum (Fig. 11). From visit 1, we obtained a noisy TS showing a feature absorption of 52 %. The TS from visit 2 clearly shows an excess absorption signal from the planetary atmosphere. For this planet, we used two Gaussian profiles to fit simultaneously the doublet and singlet lines of the He i triplet. We obtained an excess absorption of 3.500.10 for the doublet and of 0.680.12 for the singlet. The planetary signal presents a significant blue-shift (=5.60.4 km s-1). We report as our final results those from visit 2 alone (Fig. 4).
3.12 WASP-76 b
@@
The He i stellar line does not overlap with telluric lines, although the planetary trace does due to the large velocity shift during the transit (Fig 11). However, we do not detect residuals from the telluric correction. The TS of the first visit shows an absorption feature (0.550.15 %) coincident with the planetary position. The TS of the second night is a bit more noisy with an absorption feature of 0.50.1 %. Moreover, the He i line curve is mainly negative at in-transit phases (Fig. 14). When combining the spectra from both visits (Fig. 4), the final TS shows a clear absorption feature with a significant absorption of 0.480.09 % (=32 km s-1).
3.13 WASP-80 b
3.14 WASP-127 b
3.15 WASP-156 b
@@
The stellar line partially overlaps with the telluric lines, mainly in the second visit (Fig. 12), but the individual TS are mainly flat without residuals from the telluric correction. The residual map and combined TS do not present evidence of planetary absorption (Fig. 5). We derive a 3 upper limit to the He i absorption of 0.9 %.
3.16 WASP-177 b
@@
We obtained two visits with in- and out-of-transit observations in the same night, as well as out-of-transit observations in the surrounding nights. In both visits, the planetary trace does not overlap with telluric lines (Fig. 12). When analyzing visit 1 using the out-of-transit spectra from the surrounding nights, the TS shows a clear absorption-like feature. However, when using the out-of-transit spectra from the same night, this feature seems noisier. We decide to use this latter option (Fig. 12) because it minimizes any stellar variability impact. For visit 2, we found no significant differences when using the out-of-transit spectra from the same night or from one night before and after. The TS from visit 2 is mainly flat, which increases the doubts about the planetary origin of the feature from visit 1. The combined TS (Fig. 5) shows a narrow absorption due to the feature reported in the TS from visit 1. The combined residual map and TS do not provide strong evidence for a planetary absorption. We report a 3 upper limit of 1.1 %.
4 Discussion
4.1 Comparison of HPF results with literature measurements
@@
We also report the first He i triplet line spectrographic results of HAT-P-26 b, which has previously only been studied with helium narrow-band photometry by S. Vissapragada et al. (2022). We retrieve a flat TS that suggests no He i absorption, and so we set our upper limit on excess absorption at 1.2 %. However, S. Vissapragada et al. (2022) found a significant He i absorption signal with a mid-transit excess depth of 0.31 % using a 6.35 Å FWHM band-pass. We computed the planet’s helium light curve binning the TS to match the FWHM. We measure a 0.1 % excess depth, which is not consistent with the narrow-band results. The scatter present in S. Vissapragada et al. (2022)’s light curves may help to explain the discrepancy between the two observations. This discrepancy could be due to temporal planetary variability, though more observations are needed to explore this hypothesis.
@@
Additionally, we observed ten other planets, all of which have been the subject of previous spectrographic observations. For HAT-P-3 b, we set an upper limit on He i excess absorption of 1.2 %, which improves the constraint of 1.9 % set by G. Guilluy et al. (2023) using GIANO-B. Similarly, we found an upper limit of 1 % for WASP-11 b. R. Allart et al. (2023) previously set the upper limit at 1.56 % with SPIRou for this planet. While we improve previous constraints, additional observations are needed to confirm the presence of metastable helium in either planet’s atmosphere.
@@
M. Salz et al. (2016) first identified WASP-80 b as a strong candidate for atmospheric escape, particularly due to the planet’s high-energy irradiation from its host star (G. W. King et al., 2018). However, past observations by L. Fossati et al. (2022, GIANO-B), R. Allart et al. (2023, SPIRou), and A. Masson et al. (2024, SPIRou) resulted in tentative detections with upper limits of 0.7 %, 1.24 %, and 2.4 %, respectively. Our upper limit agrees with past observations and suggests that WASP-80 b does not have currently detectable helium outflow despite predictions. L. Fossati et al. (2022) considered that a sub-solar He/H ratio in combination with a high stellar wind might contribute to these non-detections. These non-detections illustrate our limited understanding of the factors that drive the development of the He i triplet.
@@
WASP-127 b has a large radius (1.31 RJup), which indicates that the planet has an extended atmosphere and is a good candidate for detecting outflowing helium (K. W. F. Lam et al., 2017). However, spectroscopic observations have only reported upper limits between 0.48-0.9 % (L. A. dos Santos et al., 2020, Gemini/Phoenix; R. Allart et al., 2023, SPIRou; A. Masson et al., 2024, SPIRou). From our HPF TS, we also put an excess absorption upper limit of 0.5 %, which is consistent with the previous results. As L. A. dos Santos et al., 2020 suggested, the metastable He state may not form due to the host star’s old age (11 Gyr, K. W. F. Lam et al., 2017) and low amount of ionizing radiation.
@@
Similarly, J. Kirk et al. (2020) identified WASP-177 b as a good candidate for the detection of He i due to its low surface gravity, large scale height, and K-type host star. However, J. Kirk et al. (2022)’s observations with Keck/NIRSPEC resulted in an upper limit of 1.25 %. This limit could be lower due to the planet’s grazing transit, and the authors noted that this tentative detection may indicate a weaker relationship between stellar high-energy activity and the formation of metastable helium. Our upper limit with HPF of 1.1 % agrees with that of J. Kirk et al. (2022).
@@
Meanwhile, we have confirmed several previous robust and tentative detections of He i for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b. We compiled those previous spectroscopic results from the ExoAtmospheres database333https://research.iac.es/proyecto/exoatmospheres/index.php, and we compare them in Figure 6.
@@
A He i triplet absorption ranging 0.35–0.70 % has been detected with several instruments for HD189733 b (M. Salz et al., 2018, CARMENES; G. Guilluy et al., 2020, GIANO-B; M. Zhang et al., 2022, Keck/NIRSPEC; A. Masson et al., 2024, SPIRou). However, we were only able to detect the signal (0.41 0.08 %) when using out-of-transit baseline contemporaneous to the in-transit observations. Otherwise, stellar variability over long timescales is greater than the planetary signal, which prevented us from initially detecting the He i triplet absorption. Therefore, our analysis for HD189733 b illustrates the limitations of the HET/HPF observing strategy for planetary atmospheric studies where stellar variability is a comparable signal or noise source to the signal.
@@
HD209458 b also displays variability in its He i signal strength. L. Nortmann et al. (2018, CARMENES) and A. Masson et al. (2024, SPIRou) set their upper limits at 0.84 % and 0.5 %, respectively. However, F. J. Alonso-Floriano et al. (2019) using CARMENES found an excess absorption of 0.91 %. Using HET/HPF, we also find an excess absorption of 0.67 %, but we note that our signal’s amplitude could be attributed to correlated noise (see Sect. 3 for further details). A variable planetary signal may have contributed to this discrepancy in detections and non-detections. We need more spectroscopic transit observations to study the behavior of the He i triplet in HD209458 b’s atmosphere.
@@
WASP-52 b has also been studied with a variety of different instruments. J. Kirk et al. (2022) and M. Fournier-Tondreau et al. (2025) reported detections of 3.44 % and 5.5 % using Keck/NIRSPEC and JWST/NIRISS, respectively. On the other hand, R. Allart et al. (2023) reported a tentative detection with an upper limit of 1.69 %. We find evidence of He i absorption, but only at a level of 1.85 0.20%, which is compatible with the upper limit reported by R. Allart et al. (2023) in the same year, but in tension with the detections. The wide spread in the reported absorptions may indicate that WASP-52 b He i signal is variable, as has been found for other exoplanets (J. P. Ninan et al., 2020; M. Zhang et al., 2022). Interestingly, the reported excess absorption levels increase with later observing dates, as seen in Figure 6. This long-term variability might be related to stellar activity W. G. Levine et al. (2024). We encourage the exoplanet community to continue observing this target in order to track its He i triplet line behavior over the next few years.
@@
As the first exoplanet with a ground-based helium detection (L. Nortmann et al., 2018), WASP-69 b is well-studied in the current literature. Past observations with a variety of instruments detect excess absorption levels ranging from 2.35–4.31 % (3.59 0.19 %, CARMENES, L. Nortmann et al., 2018; 2.21 0.46 %, SPIRou, R. Allart et al., 2023; 2.7 0.4 %, Keck/NIRSPEC, D. Tyler et al., 2024; 3.91 0.22 %, GIANO-B, G. Guilluy et al., 2024; 3.0 0.8 %, SPIRou, A. Masson et al., 2024; 3.17 0.05 %, SPIRou, R. Allart et al., 2025). We studied WASP-69b with HET/HPF for the first time. Although we retrieved a noisy TS for our first visit, it showed a tentative absorption signal. Our second visit showed a clear excess absorption of 3.5 %, consistent with previous detections in the literature.
@@
Transmission spectroscopic observations of WASP-76 b have reported ambiguous results. With CARMENES, N. Casasayas-Barris et al. (2021a) found a tentative absorption signal of 0.55 % but reported a conservative upper limit of 0.88 %. A. Masson et al. (2024) using SPIRou derived an upper limit of 0.4 % from the amplitude of the residuals near the He i triplet line. Our combined TS showed a clear absorption feature of 0.48 %, which is consistent with the previous observations. The discrepancy between results may be due to stellar variability and atmospheric changes that may have caused a decrease in the strength of the helium planetary signal, as A. Masson et al. (2024) suggested.
4.2 Helium circumstellar structures and stellar variability

### v2

@@ II.3 Transmission spectrum analysis
We analyzed the HPF spectroscopic time series using the single line transmission spectroscopy technique, which is a standard procedure that compares the in-transit spectra with the out-of-transit spectra (e.g., Wyttenbach et al., 2015; Casasayas-Barris et al., 2017). In particular, we adapted the methodology presented in Orell-Miquel et al. (2022, 2023, 2024) to the particularities of the HET following Zhang et al. (2023) and Gully-Santiago et al. (2024). Table 4 shows the planetary and stellar parameters needed to compute the transmission spectrum for our targets.
@@
Once corrected from the telluric contributions, we normalized each spectrum by its own continuum. We used muler to fit a first order Chebyshev’s polynomial to featureless spectral regions. Then, we shifted the normalized spectra into the stellar rest frame. We computed a high S/N stellar spectrum from the out-of-transit data (MasterOut). To minimize the impact of stellar variability, we analyzed each in-transit observation individually. For each transit visit, we computed a different MasterOut spectrum using only the out-of-transit spectra that are closer to the in-transit track. We usually considered all the out-of-transit spectra from a particular visit to compute the MasterOut. This option provides the most optimal results: it takes into account the behavior of the He line in short time scales, and increases the number of spectra reducing the uncertainties. On the contrary to OUT spectra, P4 spectra are usually taken far in time from the visits and the stellar He line may have varied significantly for the required precision. Moreover, P4 spectra may not be taken at optimal epochs in terms of telluric contamination. Then, we divided all the spectra (in-transit, out-of-transit, and P4) by the MasterOut to remove the stellar contribution and to obtain the residual map in the stellar rest frame. At this step, we computed the stellar light curve of the He i triplet integrating the counts within a 1.5 Å band pass centered at the doublet position. The light curves are shown in the Figures 9 and 10, and provide information of the stellar variability and the presence of large planetary tails (Zhang et al., 2023; Gully-Santiago et al., 2024).
@@
We moved the residual map into the planet rest frame accounting for the planet movement around the star. We neglected the planets’ eccentricity because it is usually poorly constrained and the impact is within the uncertainties derived for the signal shifts, except for HAT-P-17 b, HAT-P-26 b, and HD118203 b. For these three planets, we used the radvel (Fulton et al., 2018) package to compute the radial velocity contribution of the eccentricity during the transit. We computed the transmission spectrum (TS) by averaging the in-transit spectra weighted by the inverse of the squared propagated errors. We also weighted the ingress and egress spectra to account for their true contribution into the TS average (e.g., Allart et al., 2023, 2025a). In particular, we linearly interpolated from 0 at out-of-transit and 1 at full in-transit. If we obtained more than one in-transit visit, we computed the final TS from the combined residual maps. When the TS shows an absorption feature, we fitted the signal with a Gaussian profile using the Markov’s chain Monte Carlo algorithm via its python implementation emcee (Foreman-Mackey et al., 2013). If there were no clear signals in the TS, we derived conservative limits for the planetary absorption. We computed a 3 upper limit as three times the root-mean-squared value of a flat spectral region of the TS. We also estimated the upper limits in terms of equivalent width (EW) as EW = , where is the 3 upper limit in units of normalized flux, and is the average width of the He detections ( = 0.43 m; Orell-Miquel et al., 2024).
III Transmission spectroscopy results
@@
Figures 11–14 show the performance of the telluric correction and the comparison between the MasterOut and the individual in-transit spectra. For multiple visits, we also present the individual TS. The He i triplet stellar light curves are shown in Figures 9 and 10.
@@
III.1 HAT-P-3 b
The He i triplet stellar line partially overlaps with the OH lines in the only visit to the target, but the telluric emission correction is well performed (Fig. 11). The residual map and TS do not indicate the presence of a planetary absorption in our data. We found no significant differences in the TS if we only use the pre- or post-transit spectra to compute the MO, compared to the fiducial results using both sets. The TS is slightly more noisy when we add the P4 spectra into the MasterOut computation. We derive a 3 upper limit to the He i excess absorption of 1.2 %.
III.2 HAT-P-12 b
The planetary trace is overlapped by OH emission lines in the two visits (Fig. 11). While the TS from the first visit shows a broad 2 % excess absorption feature, the TS from the second visit is dominated by noise without a clear absorption signal. When combining both observations, final TS (Fig. 2) looks similar to that of the first visit. The MasterOut from the visit 1 shows a less deep He stellar line, than that from visit 2. This could explain the prominent absorption feature in visit 1. Visit 1 has pre-transit observations during the same night, but those spectra have a relatively low S/N and we obtain a less noisy TS when combining all the out-of-transit spectra for visit 1. It is evident that both transits are impacted by some OH telluric residuals, which prevent us provide a clear results. We do not consider the P4 spectra because it is affected by He airglow, and also far from both visits. Although, the S/N of the final TS would be sensitive to a He i triplet excess absorption of 1.3 %, we report a conservative upper limit of 2 % due to the telluric contamination.
III.3 HAT-P-17 b
The stellar line does not overlap with the telluric lines, but the planetary trace is in between the OH emission line and the H2O absorption line due to the RV shift caused by the planet’s eccentricity (Fig. 11). We obtained two visits which sampled the terminators of HAT-P-17 b (visit 1: ingress, visit 2: egress). Although we did not get full in-transit spectra, these observations would still be able to detect evaporation tails. Despite the telluric contamination, the TS are mainly flat without significant features (Fig. 11). However, visit 1 TS has larger uncertainties than that from visit 2, expected because we collected more spectra and with higher S/N in visit 2. There is no significant improvement when adding the P4 spectra for the MasterOut calculation. The combined TS (Fig. 2) does not present absorption features from the planet atmosphere. We derive a 3 upper limit to the He i excess absorption of 1.1 %.
III.4 HAT-P-26 b
The He i stellar line is free from tellurics, but the planetary trace falls over the blue doublet of the OH emission lines when accounting for the RV shift caused by the planet’s eccentricity (Fig. 11). Despite the telluric overlapping, the TS is mainly flat (Fig. 2) and does not suggest He i excess absorption during the transit. There is no significant improvement if we consider the few P4 spectra. The 3 upper limit to the He i excess absorption in our observations is of 1.2 %.
III.5 HD118203 b
@@
III.6 HD189733 b
We obtained three in-transit visits that sample almost all the planetary transit (visits 1, 2, and 3 in Fig. 9). The planetary trace does not overlap with the telluric lines (Fig. 12). The only observations planned as out-of-transit baseline (visit 4 in Fig. 9) are more than one month apart from the in-transit visits and the resulting TS is dominated by He i stellar line variability and without planetary absorption (see Fig. 12). Fortunately, we took two spectra as P4 observations just a few days after visit 1 (those are the two spectra at phase 0.4 colored in blue, and included as visit 1 in Figure 9). We found that the MasterOut from visit 4 and from these two P4 spectra have a 1 % discrepancy at the He stellar line. This difference is similar in magnitude to the planetary signal reported in the literature. We computed the residual map and TS shown in Fig. 3 using those two originally P4 spectra as out-of-transit and only visit 1 as in-transit. We masked some strong telluric residuals as well. The TS show a significant absorption feature at the He i triplet position (0.450.08 %). However, we emphasize that there are other features in the TS of the same amplitude.
III.7 HD209458 b
@@ III.10 WASP-52 b
The first visit is slightly contaminated by OH emission, but the TS presents an absorption at the He i triplet position (Fig. 12). We fitted the feature obtaining a blue-shifted (=73 km s-1) excess absorption of 2.40.4 % (EW = 335 mÅ). The TS of the second visit shows a clear absorption feature as well. In the second case, we fitted an excess absorption of 1.90.3 % (=13 km s-1, EW = 325 mÅ). The larger absorption and shift detected in visit 1 may be due to the slightly telluric overlap. Figure 4 shows the residual map and the TS when combining both visits. Our observations provide a clear He i triplet detection with an excess absorption of 1.850.20 %.
III.11 WASP-69 b
@@
The stellar line from both visits slightly overlap with the OH emission line, but there are no obvious telluric residuals. Although the first visit only has one in-transit spectrum with a useful S/N, the He line is clearly deeper than the MasterOut spectrum (Fig. 13). From visit 1, we obtained a noisy TS showing a feature absorption of 52 %. The TS from visit 2 clearly shows an excess absorption signal from the planetary atmosphere. For this planet, we used two Gaussian profiles to fit simultaneously the doublet and singlet lines of the He i triplet. We obtained an excess absorption of 3.500.10 % for the doublet and of 0.680.12 % for the singlet. The planetary signal presents a significant blue-shift (=5.60.4 km s-1). We report as our final results those from visit 2 alone (Fig. 4).
III.12 WASP-76 b
@@
The He i stellar line does not overlap with telluric lines, although the planetary trace does due to the large velocity shift during the transit (Fig. 13). However, we do not detect residuals from the telluric correction. The TS of the first visit shows an absorption feature (0.550.15 %) coincident with the planetary position. The TS of the second night is a bit more noisy with an absorption feature of 0.500.10 %. When combining the spectra from both visits (Fig. 4), the final TS shows a clear absorption feature with a significant absorption of 0.480.09 % (=32 km s-1). The planetary absorption is also visible in the He i line curve (Fig. 10).
III.13 WASP-80 b
The residual map and TS (Fig. 4) do not show traces of planetary absorption. The telluric lines do not overlap with the planetary trace (Fig. 13), but there are some strong telluric residuals surrounding the He i triplet position. We derive a 3 upper limit to the He i triplet excess absorption of 0.8 %.
III.14 WASP-127 b
@@
The planetary trace slightly overlaps with the blue OH doublet (Fig. 13), but the TS does not show residuals from the telluric correction. The residual map and TS are mainly flat without evidence of atmospheric absorption (Fig. 5). We derive a 3 upper limit to the He i triplet excess absorption of 0.6 %.
III.15 WASP-156 b
The stellar line partially overlaps with the telluric lines, mainly in the second visit (Fig. 13), thus the individual TS present some telluric residuals. The residual map and combined TS do not present evidence of planetary absorption (Fig. 5). We derive a 3 upper limit to the He i absorption of 0.9 %. We obtain consistent results when considering different combinations of out-of-transit spectra for the MO.
III.16 WASP-177 b
@@
We note this planet has slightly grazing transits. We obtained two visits with in- and out-of-transit observations in the same night, as well as out-of-transit observations in the surrounding nights. In both visits, the planetary trace does not overlap with telluric lines (Fig. 14). When analyzing visit 1 using the out-of-transit spectra from the surrounding nights, the TS shows a clear absorption-like feature. However, when using the out-of-transit spectra from the same night, this feature seems noisier. We decide to use this latter option (Fig. 14) because it minimizes any stellar variability impact. For visit 2, we found no significant differences when using the out-of-transit spectra from the same night or from one night before and after. The TS from visit 2 is mainly flat, which increases the doubts about the planetary origin of the feature from visit 1. The combined residual map and TS do not provide strong evidence for a planetary absorption. We report a 3 upper limit of 1.1 %.
IV Discussion
IV.1 Comparison of HPF results with literature measurements
@@
We also report the first He i triplet line spectroscopic results of HAT-P-26 b, which has previously only been studied with helium narrow-band photometry by Vissapragada et al. (2022). We retrieve a flat TS that suggests no He i absorption, and so we set our upper limit on excess absorption at 1.20 %. However, Vissapragada et al. (2022) found a significant He i absorption signal with a mid-transit excess depth of 0.31 0.10 % using a 6.35 Å FWHM band-pass. We computed the planet’s helium light curve by binning the TS to match the FWHM. We measure a 0.10 % excess depth, which is not consistent with the narrow-band results. The scatter present in Vissapragada et al. (2022)’s light curves may help to explain the discrepancy between the two observations. This discrepancy could be due to temporal planetary variability, though more observations are needed to explore this hypothesis.
@@
Additionally, we observed ten other planets, all of which have been the subject of previous spectroscopic observations. For HAT-P-3 b, we set an upper limit on He i excess absorption of 1.20 %, which improves the constraint of 1.90 % set by Guilluy et al. (2023) using GIANO-B. Similarly, we found an upper limit of 1.20 % for WASP-11 b. Allart et al. (2023) previously set the upper limit at 1.56 % with SPIRou for this planet. While we improve previous constraints, additional observations are needed to confirm the presence of metastable helium in either planet’s atmosphere.
@@
Salz et al. (2016) first identified WASP-80 b as a strong candidate for atmospheric escape, particularly due to the planet’s high-energy irradiation from its host star (King et al., 2018). However, past observations by Fossati et al. (2022, GIANO-B), Allart et al. (2023, SPIRou), and Masson et al. (2024, SPIRou) resulted in tentative detections with upper limits of 0.70 %, 1.24 %, and 2.40 %, respectively. Our upper limit agrees with past observations and suggests that WASP-80 b does not have currently detectable helium outflow despite predictions. Fossati et al. (2022) considered that a sub-solar He/H ratio in combination with a high stellar wind might contribute to these non-detections. These non-detections illustrate our limited understanding of the factors that drive the development of the He i triplet.
@@
WASP-127 b has a large radius (1.31 RJup), which indicates that the planet has an extended atmosphere and is a good candidate for detecting outflowing helium (Lam et al., 2017). However, spectroscopic observations have only reported upper limits between 0.48–0.90 % (dos Santos et al., 2020, Gemini/Phoenix; Allart et al., 2023, SPIRou; Masson et al., 2024, SPIRou). From our HET/HPF TS, we also put an excess absorption upper limit of 0.60 %, which is consistent with the previous results. As dos Santos et al., 2020 suggested, the metastable He state may not form due to the host star’s old age (11 Gyr, Lam et al., 2017) and low amount of ionizing radiation.
@@
Similarly, Kirk et al. (2020) identified WASP-177 b as a good candidate for the detection of He i due to its low surface gravity, large scale height, and K-type host star. However, Kirk et al. (2022)’s observations with Keck/NIRSPEC resulted in an upper limit of 1.25 %. This limit could be lower due to the planet’s grazing transit, and the authors noted that this tentative detection may indicate a weaker relationship between stellar high-energy activity and the formation of metastable helium. Our upper limit with HET/HPF of 1.10 % agrees with that of Kirk et al. (2022).
@@
Meanwhile, we have confirmed several previous robust and tentative detections of He i for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b. We compiled those previous spectroscopic results from the ExoAtmospheres database444https://research.iac.es/proyecto/exoatmospheres/index.php, and we compare them in Figure 6.
@@
A He i triplet absorption ranging 0.35–0.70 % has been detected with several instruments for HD189733 b (Salz et al., 2018, CARMENES; Guilluy et al., 2020, GIANO-B; Zhang et al., 2022, Keck/NIRSPEC; Allart et al., 2023; Masson et al., 2024, SPIRou). However, we were only able to detect the signal (0.45 0.08 %) when using out-of-transit baseline contemporaneous to the in-transit observations. Otherwise, stellar variability over month timescales is greater than the planetary signal, which prevented us from initially detecting the He i triplet absorption. Therefore, our analysis for HD189733 b illustrates the limitations of the HET/HPF observing strategy for planetary atmospheric studies where stellar variability is a comparable signal or noise source to the signal.
@@
HD209458 b also displays variability in its He i signal strength. Nortmann et al. (2018, CARMENES) and Masson et al. (2024, SPIRou) set their upper limits at 0.84 % and 0.50 %, respectively. However, Alonso-Floriano et al. (2019) using CARMENES found an excess absorption of 0.91 %. Using HET/HPF, we also find an excess absorption of 0.67 %, but we note that our signal’s amplitude could be attributed to correlated noise (see Sect. III for further details). A variable planetary signal may have contributed to this discrepancy in detections and non-detections. Therefore, we need more spectroscopic transit observations to study the behavior of the He i triplet in HD209458 b’s atmosphere.
@@
WASP-52 b has also been studied with a variety of different instruments. Kirk et al. (2022) and Fournier-Tondreau et al. (2025) reported detections of 3.44 % and 4.10 % using Keck/NIRSPEC and JWST/NIRISS, respectively. On the other hand, Allart et al. (2023) reported an upper limit of 1.69 %. We find evidence of He i absorption, but at a level of 1.93 0.20%, which is compatible with the upper limit reported by Allart et al. (2023) in the same year, but in tension with the detections. The wide spread in the reported absorptions may indicate that WASP-52 b He i signal is variable, as has been found for other exoplanets (Ninan et al., 2020; Zhang et al., 2022). Interestingly, the reported excess absorption levels tentatively increase with later observing dates, as seen in Figure 6. This long-term variability might be related to stellar activity (Levine et al., 2024). We encourage the exoplanet community to continue observing this target in order to track its He i triplet line behavior over the next few years.
@@
As the first exoplanet with a ground-based helium detection (Nortmann et al., 2018), WASP-69 b is well-studied in the current literature. Past observations with a variety of instruments detect excess absorption levels ranging from 2.35–4.31 % (3.59 0.19 %, CARMENES, Nortmann et al., 2018; 2.21 0.46 %, SPIRou, Allart et al., 2023; 2.70 0.40 %, Keck/NIRSPEC, Tyler et al., 2024; 3.91 0.22 %, GIANO-B, Guilluy et al., 2024; 3.00 0.80 %, SPIRou, Masson et al., 2024; 3.17 0.05 %, SPIRou, Allart et al., 2025a). We studied WASP-69b with HET/HPF for the first time. Although we retrieved a noisy TS for visit 1, it showed a tentative absorption signal. Visit 2 showed a clear excess absorption of 3.50 %, consistent with previous detections in the literature.
@@
Transmission spectroscopic observations of WASP-76 b have reported ambiguous results. With CARMENES, Casasayas-Barris et al. (2021) found a tentative absorption signal of 0.55 % but reported a conservative upper limit of 0.88 %. Masson et al. (2024) using SPIRou derived an upper limit of 0.40 % from the amplitude of the residuals near the He i triplet line. Our combined TS showed a clear absorption feature of 0.48 %, which is consistent with the previous observations. The discrepancy between results may be due to stellar variability and atmospheric changes that may have caused a decrease in the strength of the helium planetary signal, as Masson et al. (2024) suggested.
IV.2 Helium circumstellar structures and stellar variability